Night report page - 20131014

Observers: Cosimo Inserra, Ting-Wan Chen

Support team: M. Nicholl, S. Smartt

TAT on duty: Nancy Elias-Rosa, Morgan Fraser

------------------------------------------------

Subrun 3, night 2

Instruments: SOFI

WEATHER CONDITION

beginning: wind speed 2.8m/s; humidity 26%; cloudy

middle: wind speed 7.4m/s; humidity 17%; cloudy

end: wind speed 7.6m/s; humidity 18%; cloudy

CALIBRATIONS

OBSERVATIONS:

Target Type UT Filter/Grism Slit Seeing Grade (A/B/C) Comments

started at 23:30, but clouds affected focus

SN2012ca FOLLOW-UP 00:09 JHK - 1" A (telescope re-rotated, started again last two cycle, don't use the 9th frame of K-band)

HIP103361 STD 01:30 BG/RG 1. 1" A

SN2013fq FOLLOW-UP 01:38 BG/RG 1. 1" A

SN2013fs FOLLOW-UP 04:26 JHK - 1" A

SN2009ip FOLLOW-UP 04:57 HK - 1.1" A

Feige110 STD 05:51 BG/RG 1. 1" B (RG probably useless due to clouds in the direction of the target)

HIP012867 STD 06:04 BG/RG 1. 1" A (We observed this STD twice by mistake)

SN2013fc FOLLOW-UP 06:18 BG/RG 1. 1" B (RG cycle 9 out of 12)

Note: pixel scale of SOFI (large field) is 0.288 "/pixel